The measure of an angle is thirteen more than six times that of its complement. Find the measure of the angle

Answer by checkley79(3341)

Let X be the smaller angle.
let (6x+13) be the larger angle.
x+6x+13=90
7x=90-13
7x=77
x=77/7
x=11 degrees is the smaller angle.
90-11=79 degrees for the larger angle.
Proof:
11+79=90
90=90